Find the general integral of

2x(y+z2)p+y(2y+z2)q=z2

1
Expert's answer
2021-01-06T17:30:32-0500

"2x(y+z^2)p+y(2y+z^2)q=z^2\\
2x(y+z^2)\frac{\partial z}{\partial x}+y(2y+z^2)\frac{\partial z}{\partial y}=z^2\\
\frac{dx}{2x(y+z^2)}=\frac{dy}{y(2y+z^2)}=\frac{dz}{z^2}\\
\text{Multiply through by } \frac{xy}{xy}\\
y\frac{dx}{2xy^2+2xyz^2}=x\frac{dy}{2xy^2+xyz^2}=xy\frac{dz}{xyz^2}\\
\frac{ydx-xdy-xydz}{2xy^2+2xyz^2-2xy^2-xyz^2-xyz^2}=\frac{ydx-xdy-xydz}{0}..........(1)\\
ydx-xdy-xydz=0\\
\text{Divide through by } xy\\
\frac{dx}{x}-\frac{dy}{y}-dz=0\\
\text{Integrate through}\\
\ln x-\ln y-z=a\\
\ln (\frac{x}{y})-z=a"


Using "x, y, z" as multipliers. Each fraction if (1) will be

"\frac{xydx-xydy-xyzdz}{2xy^2+2xyz^2-2xy^2-xyz^2-xyz^2}=\frac{xydx-xydy-xyzdz}{0}\\
xydx-xydy-xyzdz=0\\
\text{Divude through by } xy\\
dx-dy-zdz=0\\
\text{Integrate}\\
x-y-\frac{z^2}{2}=b"


Therefore, the general solution is "f(\ln (\frac{x}{y})-z, x-y-\frac{z^2}{2})=0"